A molding machine prepares a certain kind of car spare part with a target diameter 𝜇 =




40.265 millimeters. The machine has some variability, so the standard deviation of the




diameters is 𝜎 = 0.004 millimeters. A sample of 6 spare parts is inspected each hour for




process control purposes and records are kept of the sample mean diameter. What will be




the mean and standard deviation of the numbers recorded?

1
Expert's answer
2022-04-04T16:12:35-0400

"\mu" =40.265

σ=0.004

The sampling distribution states

"\bar{x}" ="\mu"

"s=\frac{\sigma}{\sqrt{n}}"

Thus

"\bar{x}" is equal to 40.625

s="\frac{0.004}{\sqrt{6}}" =0.00163

"\frac{}{}" "\frac{}{}"